# Quadratic Mode Couplings in Rotating Black Holes and Their Detectability

### v1

@@ I Introduction
Black hole (BH) ringdown after binary BH coalescence encodes key information about the spacetime of the final BH, as well as the binary’s inspiral properties, such as the mass ratio, spins, etc [1, 2, 3, 4, 5, 6]. Theoretically, it is important to characterize various linear and non-linear signals in the ringdown stage and their connection to the inspiral parameters, as predicted by General Relativity (GR). In particular, understanding nonlinear signals not only enables direct tests of nonlinearities in GR, but also allows more theoretically meaningful ringdown fits at earlier times. This is particularly important for data analysis, as fitting the ringdown from earlier times generally leads to a higher signal-to-noise ratio (SNR).
Among various BH non-linearities [7, 8, 9, 10, 11, 4, 12, 13, 14, 15, 16, 17, 18, 19, 20], quadratic quasinormal modes (QQNMs) stand out due to their conceptual simplicity and their potential observability using current ringdown fitting techniques [21, 22, 23, 24, 25, 26, 27]. QQNMs have been identified by fitting numerical time-domain simulation data at infinity [7, 8, 9, 10, 11, 4] and on the BH horizon [14]. However, a complete and accurate characterization of the modes relies on frequency domain methods, which is the focus of this work.
Along with an earlier work [15] (hereafter Paper I), we develop two independent frequency-domain methods to handle QQNMs for generic Kerr BHs. In Paper I, we analytically reconstructed the metric and solved the frequency-domain second-order Teukolsky equation along a complex contour. Alternatively, in this paper, we use the pseudo-spectral method to numerically assemble the source term of the second-order Teukolsky equation, and then solve the equation spectrally in hyperboloidal slicing. Both methods generate consistent results for the coupling coefficient of the prograde mode $\ell=m=2,n=0$ for BH spins $\chi\in[0,0.95]$. Furthermore, we extend our discussions in Paper I by studying more observationally relevant QQNMs, and by providing a detailed classification of all possible mode-coupling channels. After examining mode coupling through different channels and symmetry conditions in (some of the) previous numerical simulations, we find consistency with the time-domain fits for mode amplitudes.
@@
Throughout this work we adopt the geometric unit by setting $G=c=1$ and we also set the remnant black hole mass $M=1$ (except for the detectability calculation). Complex conjugates are represented by overlines. We label a QNM with four indices $(l,m,n,p)$, where $p={\rm sgn}({\rm Re}(\omega))$. QNMs come in pairs. Their frequency values are related via
@@ II QQNMs with hyperboloidal slicing
QNMs are known to diverge at spatial infinity and at the bifurcate horizon. While this divergence poses no physical issue because they do not lie in the region where the QNMs are a good approximation, it can cause numerical challenges. In Paper I, we deal with the divergence by using the complex plane approach. Here, we adopt an alternative method — hyperboloidal time slicing — which avoids the problematic regions. This enables direct numerical treatments of the second-order Teukolsky equation [31]
@@
thus $h_{ab}^{(2)}$ satisfies the linearized Einstein equations with $-G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]/8\pi$ acting as the effective stress-energy tensor at this order. After computing $G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]$ from the linearized metric $h_{ab}^{(1)}$, we convert it to the source $S_{4}^{(2)}$ for $\Psi_{4}^{(2)}[h_{ab}^{(2)}]$ in Eq. (2), as well as the source for $\Psi_{0}^{(2)}[h_{ab}^{(2)}]$.
@@ III Classification of mode couplings
With this notation, several key observations can be made. First, the QQNM appears only in the $(m_{1}+m_{2})$ GW harmonics due to the azimuthal angular selection rule. Second, the notation exhausts all possible QQNMs. Third, different pairs of parent modes (up to exchange symmetry $1\leftrightarrow 2$) generally produce distinct frequencies. An exception occurs for Schwarzschild BHs, where spherical symmetry makes QNM frequencies independent of $m$. Nonetheless, different $(m_{1},m_{2})$ pairs indicate distinct coupling channels. The total QQNM amplitude in a GW harmonic is the sum of all possible pairs $(m_{1},m_{2})$ with $m_{1}+m_{2}$ fixed.
@@ V Forecast on detectability
In the following, we focus on CE [37] and LISA [38]. For simplicity, we average the SNRs over all angles, which yields the orthogonal condition between two antenna patterns $F_{+,\times}$ for CE:
@@
The QQNM $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$ was also identified in numerical relativity waveforms. In Fig. 2a, we present its SNR dependence on the mass ratio $q$ and the effective inspiral spin $\chi_{\rm eff}$. The optimal binary configuration is $q\sim 2.75$ with maximally positive $\chi_{\rm eff}$. By setting the total binary mass $M_{\rm tot}$ to $100M_{\odot}$, and the luminosity distance $D_{L}$ to 410 Mpc, the SNR detected by CE can exceed 8 for some parameter space. With LISA, the SNR dependence on $q$ and $\chi_{\rm eff}$ remains the same. Assuming $M_{\rm tot}=8\times 10^{6}M_{\odot}$ and a redshift of 1, we find its SNR ranges from 16 to 653. The corresponding plot is available in the Supplementary Material.
@@ Acknowledgements.
Acknowledgments.— We thank Hengrui Zhu for sharing the data in Fig. 1. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. N.K. is supported by the Natural Science and Engineering Council of Canada.

### v2

@@ I Introduction
Black hole (BH) ringdown after binary BH coalescence encodes key information about the spacetime of the final BH [1, 2, 3], as well as the binary’s inspiral properties, such as the mass ratio, spins, etc [4, 5, 6, 7, 8, 9]. Theoretically, it is important to characterize various linear and non-linear signals in the ringdown stage and their connection to the inspiral parameters, as predicted by General Relativity (GR). In particular, understanding nonlinear signals not only enables direct tests of nonlinearities in GR, but also allows more theoretically meaningful ringdown fits at earlier times. This is particularly important for data analysis, as fitting the ringdown from earlier times generally leads to a higher signal-to-noise ratio (SNR).
Among various BH non-linearities [10, 11, 12, 13, 14, 7, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], quadratic quasinormal modes (QQNMs) stand out due to their conceptual simplicity and their potential observability using current ringdown fitting techniques [27, 28, 29, 30, 31, 32, 33]. QQNMs have been identified by fitting numerical time-domain simulation data at infinity [10, 11, 12, 13, 14, 7] and on the BH horizon [17]. However, a complete and accurate characterization of the modes relies on frequency domain methods, which is the focus of this work.
Along with an earlier work [18] (hereafter Paper I), we develop two independent frequency-domain methods to handle QQNMs for generic Kerr BHs. In Paper I, we analytically reconstructed the metric and solved the frequency-domain second-order Campanelli–Lousto-Teukolsky equation [34] along a complex contour. Alternatively, in this paper, we adopt the reduced second-order Teukolsky equation [35, 36] and use the pseudo-spectral method to numerically assemble its source term. Subsequently, we solve the equation spectrally in hyperboloidal slicing [37, 38, 39]. Both methods generate consistent results for the coupling coefficient of the prograde mode $\ell=m=2,n=0$ for BH spins $\chi\in[0,0.95]$. Furthermore, we extend our discussions in Paper I by studying more observationally relevant QQNMs, and by providing a detailed classification of all possible mode-coupling channels. After examining mode coupling through different channels and symmetry conditions in (some of the) previous numerical simulations, we find consistency with the time-domain fits for mode amplitudes.
@@
Throughout this work we adopt the geometric unit by setting $G=c=1$. Complex conjugates are represented by overlines. We label a QNM with four indices $(l,m,n,p)$, where $p={\rm sgn}({\rm Re}(\omega))$. QNMs come in pairs. Their frequency values are related via
@@ II QQNMs with hyperboloidal slicing
QNMs are known to diverge at spatial infinity and at the bifurcate horizon. While this divergence poses no physical issue because they do not lie in the region where the QNMs are a good approximation, it can cause numerical challenges. In Paper I, we deal with the divergence by using the complex plane approach. Here, we adopt an alternative method — hyperboloidal time slicing — which avoids the problematic regions. This enables direct numerical treatments of the reduced second-order Teukolsky equation [35, 36]
@@
thus $h_{ab}^{(2)}$ satisfies the linearized Einstein equations with $-G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]/8\pi$ acting as the effective stress-energy tensor at this order. After computing $G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]$ from the linearized metric $h_{ab}^{(1)}$, we convert it to the source $S_{4}^{(2)}$ for $\Psi_{4}^{(2)}[h_{ab}^{(2)}]$ in Eq. (2). Note that while Eq. (5) only provides the source for the part of $\Psi_{4}$ that depends linearly on $h_{ab}^{(2)}$, in ORG quadratic contribution to $\Psi_{4}$ from $h_{ab}^{(1)}$ vanishes.
@@ III Classification of mode couplings
With this notation, several key observations can be made. First, the QQNM appears only in the $(m_{1}+m_{2})$ GW harmonics due to the azimuthal angular selection rule. Second, the notation exhausts all possible QQNMs. Third, different pairs of modes (up to exchange symmetry $1\leftrightarrow 2$) generally produce distinct frequencies. An exception occurs for Schwarzschild BHs, where spherical symmetry makes QNM frequencies independent of $m$. Nonetheless, different $(m_{1},m_{2})$ pairs indicate distinct coupling channels. The total QQNM amplitude in a GW harmonic is the sum of all possible pairs $(m_{1},m_{2})$ with $m_{1}+m_{2}$ fixed.
@@ V Forecast on detectability
In the following, we focus on a 40 km CE [48] and LISA [49]. For simplicity, we average the SNRs over all angles, which yields the orthogonal condition between two antenna patterns $F_{+,\times}$ for CE:
@@
The QQNM $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$ was also identified in numerical relativity waveforms. In Fig. 2a, we present its SNR dependence on the mass ratio $q$ and the effective inspiral spin $\chi_{\rm eff}$. For simplicity, we assume both component spins $\chi_{1,2}$ are equal, namely $\chi_{1}=\chi_{2}=\chi_{\rm eff}$. The optimal binary configuration is $q\sim 2.75$ with maximally positive $\chi_{\rm eff}$. By setting the total binary mass $M_{\rm tot}$ to $100M_{\odot}$, and the luminosity distance $D_{L}$ to 410 Mpc, the SNR detected by CE can exceed 8 for some parameter space. With LISA, the SNR dependence on $q$ and $\chi_{\rm eff}$ remains the same. Assuming $M_{\rm tot}=8\times 10^{6}M_{\odot}$ and a redshift of 1, we find its SNR ranges from 16 to 653. The corresponding plot is available in the Supplementary Material.
@@ Acknowledgements.
VII Acknowledgments
We thank Hengrui Zhu for sharing the data in Fig. 1. We thank Andrew Spiers for pointing out the difference between the Campanelli–Lousto–Teukolsky equation and the reduced second-order Teukolsky equation. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. N.K. is supported by the Natural Science and Engineering Council of Canada.